OpenStudy (i_ask_alot):

Help A$AP!!! A neighborhood has 32 one-story homes and 24 two-story homes. What is the ratio of two-story homes to one-story homes in simplest form?

OpenStudy (i_ask_alot):

Is it 3/4?

OpenStudy (anonymous):

we need the answers so we can help

OpenStudy (anonymous):

24/32 is the ratio. Now let's simplify it: 24/8=3...32/8=4 The simplest ratio is 3/4

OpenStudy (tkhunny):

Ratio -- Fraction -- Division Problem Set it up and simplify it. Don't provide answer choices, just get it right.

OpenStudy (anonymous):

yeah its 3:4

OpenStudy (i_ask_alot):

Bye

OpenStudy (tkhunny):

Next time, show YOUR work.